OpenStudy (anonymous):

George makes large cookies. These cookies are 18 inches in diameter. How much icing will he need to cover the top of each cookie?

OpenStudy (chriss):

He will need 18pi square inches of frosting per cookie.

OpenStudy (anonymous):

You are good at this :)

OpenStudy (anonymous):

This answer was wrong.

OpenStudy (anonymous):

Thanks for the effort though!

OpenStudy (chriss):

right... I was still thinking circumference... I should have used the formula pir^2 r=1/2d so r=9 r^2=81 A=81pi so 81pi square inches Sorry for the initial error.